Find two numbers given that their sum is 48 and the smaller number is equal to 1/5 of the larger number.
The answer at the back of the book is 40,8.

Answer by mananth(16946)

Find two numbers given that their sum is 48 and the smaller number is equal to 1/5 of the larger number.

let one number be x
other be y (x>y)
x+y =48...................(1)
y= (1/5) x
5y=x
x+y =48
substitute x
5y+y =48
6y=48
y =8
Now x+y =48
x+8=48
x=40
40, 8